Find equilibrium Price and Quantity and Price Elasticity of Demand from the following demand and supply functions

         i.       Q=10-2P                     Q=5+3P

         ii.      Qd= 200-40P                Qs= -40+80P


1
Expert's answer
2021-03-23T08:23:13-0400

i) The market equilibrium occurs when "Q_d=Q_s":


"10-2P_E=5+3P_E,""5=5P_E,""P_E=\$1."

Then, we can find the equilibrium quantity:


"Q_E=10-2\cdot1=8."

Let's write the inverse-demand function:


"P=5-0.5Q."

The price elasticity of demand can be found as follows:


"E_d=(\dfrac{-1}{slope})\dfrac{P}{Q_d},""E_d=(\dfrac{-1}{-0.5})\cdot\dfrac{1}{8}=0.25."

ii) The market equilibrium occurs when "Q_d=Q_s":


"200-40P_E=-40+80P_E,""240=120P_E,""P_E=\$2."

Then, we can find the equilibrium quantity:


"Q_E=200-40\cdot2=120."

Let's write the inverse-demand function:


"P=5-0.025Q."

The price elasticity of demand can be found as follows:


"E_d=(\dfrac{-1}{slope})\dfrac{P}{Q_d},""E_d=(\dfrac{-1}{-0.025})\cdot\dfrac{2}{120}=0.67."